A moving coil instruments gives a full scale deflection with a potential difference of 70mV and a current of 20mA calculate the value of the shunt required to give a range of 0-50Amps.
1
Expert's answer
2020-02-06T09:27:27-0500

As per the given question,

the potential difference with the full scale deflection=70mV

The current reading with the full scale deflection=20mA

Resistance of the shunt=?

Range of the current =0-50A

So, "G=\dfrac{V}{i}=\dfrac{70}{20}=3.5\Omega"

We know that, "S=\dfrac{I_G \times G}{I-I_G}"

"S=\dfrac{3.5\times20\times10^{-3}}{50-20\times 10^{-3}}=\dfrac{0.07}{50-0.02}"

"S=\dfrac{0.07}{49.08}=0.00142\Omega"

Or, S="1.42\times 10^{-3}\Omega"